# Efficient Monte-Carlo sampling of metastable systems using non-local collective variable updates

### v2

@@ H.4 Dimer in a solvent
In addition to Fig.˜8, we ran the algorithm on the dimer model for additional values of $\alpha_{1}$. The associated inverse mode-jump cost is shown in Fig.˜17. Apart from the deterministic algorithm ($\alpha_{1}=0$), almost no mode switches were observed within the computational budget.
Figure 17: Mode jump cost for the Dimer model from Section˜III.3 for different parameters, estimated from running 20 MCMC chains over 5,000 iterations. The inverse dimensionless velocity $1/\widetilde{v}$ corresponds to the number of steps for a steered schedule over a unit distance. White space indicates that at least one chain never switched mode during observation time. The left-most panel shows the same results as Fig.˜8 (with a different color scale).

### v3

@@ H.4 Dimer in a solvent
In addition to Fig.˜9, we ran the algorithm on the dimer model for additional values of $\alpha_{1}$. The associated inverse mode-jump cost is shown in Fig.˜18. Apart from the deterministic algorithm ($\alpha_{1}=0$), almost no mode switches were observed within the computational budget.
Figure 18: Mode jump cost for the Dimer model from Section˜III.3 for different parameters, estimated from running 20 MCMC chains over 5,000 iterations. The inverse dimensionless velocity $1/\widetilde{v}$ corresponds to the number of steps for a steered schedule over a unit distance. White space indicates that at least one chain never switched mode during observation time. The left-most panel shows the same results as Fig.˜9 (with a different color scale).